An electric kettle contains 1.5 litres of water at 20°C. Find;
a) the heat required to bring it to boiing point; [Ans.: 504 000 J]
b) the additional heat necessary to boil half the water away, assuming that all the heat from the element goes into the water. [Ans.: 1 695 000 J]
c) If the element is rated 1 000 W, how long does it take the water to come to the boil? [Ans.: 8 min 24s.]

Answer by math_helper(2461)


It takes 4200J to raise 1kg of water by 1°C
a) To boil, you must raise the temperature to 100°C (1 atmosphere of pressure assumed). That's a change of 100-20 = 80°C
1.5l has a mass of 1500g or 1.5kg
80*1.5*4200 = 504000J
Same calculation including unit analysis:
80°C*1.5kg*4200J/kg/°C = 504000J

b) It takes 2260J/g of heat to turn water at 100°C into steam.
One half of the water in this case is 0.5*1.5kg = 0.75kg
2260J/g = 2260000J/kg * 0.75kg = 1695000J

c) One watt is one joule per second (1W = 1J/s) Watts are units for POWER (work per unit time) while Joules are units of ENERGY or WORK.
1000W ==> 1000J/s * t work can be done in t seconds
504000J = 1000J/s * t ==> t = 504sec ( = 8min 24sec )